# Neural Spelling: A Spell-Based BCI System for Language Neural Decoding

### v1

@@ V-E Stage 1: Neural Letter Classifier
V-E1 Model Performance Comparison
In this section, we evaluate the efficacy of different neural network architectures for EEG-based encoding by comparing a CNN with both Long Short-Term Memory (LSTM) networks and Transformer-based models, each equipped with and without CL module. The comparative analysis is rooted in statistical testing and performance metrics across three classification accuracies: Top 1, Top 3, and Top 5.
TABLE II: Main Decoding Results
@@ V-E1 Model Performance Comparison
TABLE III: Ablation on the impact of space separation between words during neural spelling
@@
TABLE IV: Ablation Study on the impact of letter classification sampling space
@@
The CNN augmented with the CL module significantly outperformed the other models. Specifically, the CNN w/ CL achieved a Top 1 accuracy of $33.10\%\pm 11.54\%$, a Top 3 accuracy of $57.77\%\pm 12.89\%$, and a Top 5 accuracy of $71.46\%\pm 11.11\%$. Statistical analyses, including pairwise t-tests, revealed that the CNN with CL model was superior to the LSTM w/ CL (t(334) = 9.505, $\textit{p}~{}<~{}0.001$) and the Transformer w/ CL (t(334) = 12.028, $\textit{p}~{}<~{}0.001$). All tests are corrected by Benjamini-Hochberg False Discovery Rate (FDR_BH) to avoid multi-comparison issues, ensuring robustness in reporting statistically significant differences.
Fig. 3 delineates the comparative accuracy of these models across the Top 1, Top 3, and Top 5 metrics, where asterisks denote statistical significance at the $\textit{p}~{}<~{}0.001$ level.
@@ V-E2 Model Performance Across Different ROIs and Frequencies
This section investigates the impact of different ROIs and frequency bands on the Top 1 accuracy of a CNN model augmented with CL. The analysis focuses on the PSD extracted from specific bands and ROIs to determine their relative importance in performance metrics.
@@ V-F2 Sentence Decoding Performance
Table II presents the results of sentence decoding from the encoder, utilizing the Top-k predicted distribution as input for the LLM decoder (where $k=3$). We evaluate the decoding performance across two different model sizes of the BART model [45], namely ‘large‘ and ‘base‘, with the trainable parameters equal to 406M and 140M, respectively. For both model sizes, our proposed curriculum learning method significantly enhances decoding performance across all evaluated metrics. Specifically, using the pretrained BART-large model, our method achieves a BLEU-4 score of $44.42\%\pm 2.2\%$ and a ROUGE-L score of $63.71\%\pm 0.86\%$.
@@
Fig. 5(B) revealing a positive linear relationship between the decoder’s performance on the sentence decoding task and the neural letter classifier’s performance. It shows that LLMs after neural letter classifier can significantly improve in word level.
@@ V-F3 Visualization of Sentence Decoding Results
TABLE V: Main Decoding Results
Type Content
Classifier Output rn thp unltrgfving wflyernesf kl the amaztn fafnforelt i sroup os rplgrfchprs lgd bu da rtpria srnvhpz uncowgr r grydgn cicilrzaorkn dggp wrohfn thg kunsle
Ground Truth In the unforgiving wilderness of the Amazon rainforest, a group of researchers led by Dr. Amelia Sanchez uncover a hidden civilization deep within the jungle.
LM Decoder In the unforgiving winds of the Amazon rainforest, a group of survivors led by Dr. Anya Singh uncover a hidden civilization that harnessed the power of nature’s healing properties.
Classifier Output al hef creitikns bpuowg fncrgrsingry xtoyfrf fhe frups thg gthfcrl dfrgtma of trnfxulrtrnj xpooref orlts ind the ctnsequencel ol blyfrrng zhg rfnpf bgtwepn rparity rnd srbrication
Ground Truth As her creations become increasingly powerful, she faces the ethical dilemma of manipulating people’s emotions and the consequences of blurring the lines between art and reality.
LM Decoder As her creations become increasingly popular, she faces the ethical dilemma of manipulating people’s pasts and the consequences of blurring the lines between reality and fabrication.
Classifier Output tfiver r cybtrg detpctivp mith the ibilrty tt rewind trmg fnceszfgazgs tempoaar crrtef rn r futurfszrv tetroooril
Ground Truth Oliver, a cyborg detective with the ability to rewind time, investigates temporal crimes in a futuristic metropolis.
LM Decoder Thomas, a cybernetic detective with the ability to rewind time, investigates temporal anomalies in a futuristic metropolis.
Classifier Output as ag chrfef yown rogue trmp trrvglgrl and ynaavpff uknsxifauiel ag unvocprf i lakcking zruth rbout hff omn crpaoitn rnd fts oiel oo thg very frbrru ks tfte skauing hrt zt uhtklg setwgen lgavrnj duszice tr embrrcrng hrl taue oufxole even rs io mgrns rloering zhe ctyrse of hiltkay
Ground Truth As he chases down rogue time travelers and witnesses unimaginable events, he uncovers a shocking truth about his own creation and its potential to shape the future of time, forcing him to choose between preserving history or allowing it to evolve naturally.
LM Decoder As he chases down rogue time travelers and unravels conspiracies, he uncovers a shocking truth about his own creation and its ties to the very fabric of time, forcing him to choose between serving justice or embracing his true purpose, even if it means altering the course of history.
Classifier Output fn a xoltaotcrlyotrc waftefiny krf i rgsoyruesur lvavenger ltumbles uxtn r soagkzten rfvhrvp cknzainrng loabfkken knomlpdgg rbtut thp wkrld bestap tae ctrrrolg
Ground Truth In a post-apocalyptic wasteland, Kai, a resourceful scavenger, stumbles upon a forgotten archive containing forbidden knowledge about the world before the collapse.
LM Decoder In a post-apocalyptic world, Kai, a resourceful scavenger, stumbles upon a forgotten archive containing forgotten knowledge about a legendary artifact.
Classifier Output kn the dusty tarfls ks thg grrnd vrnutn srgp r nowrdrv artilt wioa a paratbrush rf hea gurdg lharpl ftkrfgs wrza zeohya r wffe old ckykte wgo underftrnyl the rgyohmf tl thg kesgrt
Ground Truth On the dusty trails of the Grand Canyon, Sage, a nomadic artist with a paintbrush as her guide, shares stories with Zephyr, a wise old coyote who understands the rhythms of the desert.
LM Decoder On the dusty plains of the Grand Canyon, Sage, a nomadic artist with a brush as her guide, shares stories with Whispering, a wise old Cherokee woman who understands the language of the land.
Note: All letters in the displayed results have been automatically converted from uppercase to lowercase for visualization.
For qualitative analysis, we visualize the decoding results of our method in Table V. The input samples were taken from subject P11, which exhibits the highest performance in letter classification. Notably, some phrases from the initial decoding are already remarkably similar to the ground truth. For instance, ’rn thp unltrgfving…’ closely approximates ’in the unforgiving…’. Our LLM decoder capitalizes on the word-transitional distribution, facilitating improvements in sentence grammar and word-level corrections. As a result, the decoded sentences closely align with the ground truth, demonstrating the efficacy of our approach in refining linguistic outputs.
@@ V-G1 Ablation on the effect of letter sampling range
In our analysis, we observe a decline in performance across all metrics when sampling from a larger prediction range. This outcome illustrates a trade-off between letter-level prediction accuracy and input diversity. To balance these factors effectively, we opt to utilize the top-3 prediction results for each letter to sample the letters spelled by the user, as detailed in Table IV. This approach optimizes our model’s accuracy while maintaining a reasonable level of diversity in the input data.
@@ V-G2 Ablation on the Effect of Space Between Words During Neural Spelling
Results are presented in Table III. In conclusion, while incorporating spaces between words introduces an additional step for the user, it aids the model in more effectively denoising the input. However, the overall results demonstrate that our model is capable of denoising neural spelling text effectively even when spaces are not included in the input. This underscores the robustness of our approach in handling continuous streams of characters without explicit segmentation.
@@ VI Discussion
This study presents a Curriculum-based Neural Spelling Framework comprising a Convolutional EEG Encoder and a CL module, demonstrating the feasibility of recognizing handwritten letters using non-invasive technologies. Furthermore, the system highlights the potential for real-world applications by integrating LLMs to generate free-form sentences based on spelling-based designs.
@@ VI-A Neurophysiological Correlates of Letter Recognition
Figure 8: UMAP visualization of letter-specific neural representations. (A): EEG PSD features with averaged baseline removed to enhance visualization clarity. (B): Handwriting trajectories showing distinct clusters for letters TWY, BFE, ZK, and OC, which suggest similar neurocognitive processing pathways and minor differences in movement patterns. (C): Significant differences between letter pairs A and Z, B and L, B and J, and O and Z, assessed via paired t-tests and corrected for multiple comparisons using the FDR_BH method (colors indicate $\textit{p}<0.05$), highlighting the neurophysiological distinctions in letter processing.
Handwriting involves complex neural coordination, engaging multiple brain regions. Figure 8 illustrates how specific letters correspond to neural activation (A) and handwriting trajectory (B). Some letters, such as TWY, BFE, ZK, and OC, show similar patterns in both neural and trajectory patterns. The small distance of these letters, suggests a shared representation in neural processing spaces.
@@
Additionally, the sensorimotor cortex’s involvement aligns with its role in governing motor control and sensory processing, fundamental to handwriting [60]. The broad gamma activity across these regions suggests a high-level synchronization of cortical activity, facilitating coherent cognitive representations necessary for complex tasks like letter recognition and differentiation [61, 62, 63].
@@ VI-B Inter-Trial Coherence in Delta Band During Handwriting Tasks
Studies have indicated the involvement of Delta and Theta bands in motor tasks and cognitive processing. For instance, increased ITC in these bands has been associated with task switching and inhibitory control in BCI paradigms [66]. Similar patterns of coherence in Delta-Theta ranges have been observed during motor performance improvements in stroke rehabilitation, reflecting neural reorganization and recovery [67]. The periodic auditory stimulation research shows a synchronization peak at 2 Hz in the Delta band, linking sensory-motor time coupling, which may parallel the neural dynamics observed during the rhythmic movements in handwriting [68].
@@ VII Conclusions
This study introduces a Curriculum-based Neural Spelling Framework (CNS) that leverages the advanced capabilities of GenAI to enhance spell-based neural language decoding tasks. Our approach is distinct in integrating a CNN with a curriculum-driven LLM, promoting an innovative hybrid method in the domain of BCIs. The framework’s effectiveness is demonstrated through its application to EEG-based handwriting of all 26 letters, a novel endeavor in the field. The CNS framework notably achieves exemplary top-k accuracy across all subjects, underscoring the robustness of the EEG encoding model. Furthermore, our curriculum supervised fine-tuning method significantly advances the state of the art by enabling the LLM to effectively learn subject-specific letter transition patterns. This methodological innovation not only enhances sentence synthesis quality but also sets a new benchmark for assistive communication technologies. By seamlessly merging non-invasive EEG with GenAI, this study not only addresses the immediate needs of individuals with diverse physical abilities but also sets the stage for future explorations in sophisticated, accessible communication solutions. As we continue to refine these integrations, the potential to expand the capabilities of BCIs and improve the quality of life for users worldwide remains vast and inspiring.
@@ VIII Limitations
Despite its notable strengths, the proposed framework exhibits several limitations that warrant further investigation. Firstly, there is a need to collect and validate online sentence data to ascertain the framework’s efficacy in real-world applications. Secondly, the current task framework is primarily based on within-subject analyses, limiting its generalizability across different subjects. The capacity for cross-subject transfer learning remains underexplored and requires significant enhancement to ensure broader applicability. Thirdly, the dataset size for individual subjects is relatively small, which constrains the training of a more extensive and robust brain encoder necessary for a high-performing letter classifier. Future work should focus on collecting more comprehensive data sets and developing a robust, transferable model that can operate effectively across subjects in real-world, online applications.

### v2

@@ V-E Stage 1: Neural Letter Classifier
TABLE III: Main Decoding Results
@@
TABLE IV: Ablation on the impact of space separation between words during neural spelling
@@
TABLE V: Ablation Study on the Impact of Letter Classification Sampling Space
@@
V-E1 Model Performance Comparison
We evaluated the performance of different neural network architectures for EEG-based letter decoding, including LSTM, transformer-based, EEG-Deformer [15], Conformer [65], EEGFormer [70], and TCNet [32], with and without ConL. The comparison focuses on Top-1, Top-3, and Top-5 classification accuracy and is summarized in Table II.
Across all evaluated models, the proposed CNN-based architecture consistently achieved the highest decoding accuracy under the current experimental setting. In particular, the CNN with ConL attained a Top-1 accuracy of $33.10\%\pm 11.54\%$, a Top-3 accuracy of $57.77\%\pm 12.89\%$, and a Top-5 accuracy of $71.46\%\pm 11.11\%$, outperforming all recurrent and transformer-based baselines. Paired t-tests across subject-sessions indicate that these improvements are statistically significant in most comparisons, with multiple-comparison correction applied using the Benjamini–Hochberg procedure.
Confusion matrix analysis.
Fig. 5(A) illustrates the confusion matrix of the averaged conditional probability $P(\text{predicted}\mid\text{true})$ for 26 handwritten letters. Overall decoding performance varies substantially across letters. Letter I achieves the highest classification accuracy (33.12%), whereas letter R exhibits the lowest accuracy (15.00%), indicating pronounced letter-specific decoding difficulty.
Certain letters show consistently higher confusion rates. In particular, letter E demonstrates relatively poor performance (23.13%) and is frequently misclassified as visually or motorically similar letters, including D, J, G, N, B, H, R, L, S, and X. This widespread confusion suggests that E shares common stroke components and motor primitives with multiple letter classes, making it difficult to disentangle under non-invasive EEG constraints.
Fig. 5(B) further summarizes the most frequently confused letter pairs. Among them, the M–N pair exhibits the highest mutual confusion probability, which can be attributed to their highly similar handwriting trajectories characterized by repeated vertical strokes and comparable temporal motor patterns. Overall, these results indicate that misclassifications are highly structured rather than random, reflecting intrinsic similarities in neural and motor representations rather than decoder instability.
Figure 5: (A): Confusion matrix of letter decoding performance with hierarchical clustering. The matrix shows the averaged conditional probability $P(\text{predicted}\mid\text{true})$ (colored $P>0.1$) for 26 handwritten letters, reordered using hierarchical clustering based on confusion profiles. (B): Top-10 Most Frequently Confused Letter Pairs.
@@ V-E2 Model Performance Across Different ROIs and Frequencies
This section investigates the impact of different ROIs and frequency bands on the Top 1 accuracy of a CNN model augmented with ConL. The analysis focuses on the PSD extracted from specific bands and ROIs to determine their relative importance in performance metrics.
@@
Informative EEG Feature Analysis. To explicitly investigate which EEG features are most informative for letter decoding, we analyzed decoding performance across both spectral and spatial dimensions. As shown in Fig. 3, frequency-band analysis reveals that high-frequency components, particularly the gamma band (30–70 Hz), contribute most strongly to Top-1 decoding accuracy. In contrast, lower-frequency bands (delta and theta) show relatively limited discriminative power for letter classification. Spatially, electrodes over the prefrontal and frontal cortices yield the highest decoding performance, followed by temporal and sensorimotor regions. Using all ROIs jointly achieves the best overall accuracy, indicating that neural letter representations are distributed across multiple cortical areas rather than localized to a single region. Together, these results suggest that informative features for non-invasive neural spelling are largely associated with high-frequency spectral activity in frontal–prefrontal networks associated with motor planning and language-related processing.
@@ V-F2 Sentence Decoding Performance
Table III presents the results of sentence decoding from the encoder, utilizing the Top-k predicted distribution as input for the LLM decoder (where $k=3$). We evaluate the decoding performance across two different model sizes of the BART model [39], namely `large` and `base`, with the trainable parameters equal to 406M and 140M, respectively. For both model sizes, our proposed curriculum learning method significantly enhances decoding performance across all evaluated metrics. Specifically, using the pretrained BART-large model, our method achieves a BLEU-4 score of $44.42\%\pm 2.2\%$ and a ROUGE-L score of $63.71\%\pm 0.86\%$.
@@
Fig. 4(B) revealing a positive linear relationship between the decoder's performance on the sentence decoding task and the neural letter classifier's performance. It shows that LLMs after neural letter classifier can significantly improve in word level.
@@ V-F3 Visualization of Sentence Decoding Results
For qualitative analysis, we visualize the decoding results of our method in Supplemental Table I. The input samples were taken from subject P11, who exhibited the highest letter-level classification performance in our dataset. Notably, some phrases from the initial neural decoding are already partially informative and resemble the target sentence structure. For example, the fragment `rn thp unltrgfving…' bears clear similarity to the ground-truth phrase `in the unforgiving…'.
The LLM decoder then integrates these noisy letter-level predictions with learned word-transition statistics and linguistic constraints, enabling improved grammatical structure and word-level coherence. Importantly, the resulting sentences are not intended to be exact verbatim reconstructions of the ground truth. Instead, they represent semantically plausible and communicatively usable outputs that balance neural evidence with strong language priors.
As illustrated in Supplemental Table I, this process may introduce creative variations, particularly when the neural input is highly uncertain, reflecting a trade-off between symbol-level fidelity and language-level coherence. These qualitative results demonstrate the effectiveness of the proposed framework in refining linguistic outputs under noisy neural decoding conditions, while preserving consistency with the underlying neural evidence.
@@ V-G1 Ablation on the effect of letter sampling range
In our analysis, we observe a decline in performance across all metrics when sampling from a larger prediction range. This outcome illustrates a trade-off between letter-level prediction accuracy and input diversity. To balance these factors effectively, we opt to utilize the top-3 prediction results for each letter to sample the letters spelled by the user, as detailed in Table V. This approach optimizes our model's accuracy while maintaining a reasonable level of diversity in the input data.
@@ V-G2 Ablation on the Effect of Space Between Words During Neural Spelling
Results are presented in Table IV. In conclusion, while incorporating spaces between words introduces an additional step for the user, it aids the model in more effectively denoising the input. However, the overall results demonstrate that our model is capable of denoising neural spelling text effectively even when spaces are not included in the input. This underscores the robustness of our approach in handling continuous streams of characters without explicit segmentation.
@@ VI Discussion
This study presents a Curriculum-based Neural Spelling Framework comprising a Convolutional EEG Encoder and a ConL module, demonstrating the feasibility of recognizing handwritten letters using non-invasive technologies. Furthermore, the system highlights the potential for real-world applications by integrating LLMs to generate free-form sentences based on spelling-based designs.
@@ VI-A Neurophysiological Correlates of Letter Recognition
Figure 8: UMAP visualization of letter-specific neural representations. (A): EEG PSD features with averaged baseline removed to enhance visualization clarity. (B): Handwriting trajectories showing distinct clusters for letters TWY, BFE, ZK, and OC, which suggest similar neurocognitive processing pathways and minor differences in movement patterns. (C): Significant differences between letter pairs A and Z, B and L, B and J, and O and Z, assessed via paired t-tests and corrected for multiple comparisons using the FDR_BH method (colors indicate $\textit{p}<0.05$), highlighting the neurophysiological distinctions in letter processing. UMAP embeddings were computed and visualized on a per-subject basis to account for inter-individual EEG variability.
From a feature perspective, our results indicate that letter decoding is strongly associated with spectral-spatial EEG patterns rather than transient time-domain signals. In particular, gamma-band activity over frontal and prefrontal regions exhibits strong discriminative power across letters. This pattern is consistent with prior literature linking gamma activity to fine motor execution and higher-order cognitive processing.
In contrast, alpha-band activity shows letter-dependent modulation mainly in parietal and central regions, which is consistent with visuomotor integration and trajectory planning during handwriting. Lower-frequency oscillations, including delta and theta bands, demonstrate increased inter-trial coherence during movement initiation but contribute less directly to class separability. These findings are consistent with a potential functional differentiation in which high-frequency spectral features are more strongly associated with letter-specific decoding performance, while low-frequency dynamics appear more related to task coordination and motor timing.
The qualitative analysis of the Event-Related Potential (ERP) results is presented in Supplemental Section II.A.
Handwriting involves complex neural coordination, engaging multiple brain regions. Fig.  8 illustrates how specific letters correspond to neural activation (A) and handwriting trajectory (B). Some letters, such as TWY, BFE, ZK, and OC, show similar patterns in both neural and trajectory patterns. The small distance of these letters, suggests a shared representation in neural processing spaces.
@@
Additionally, the sensorimotor cortex's involvement aligns with its role in governing motor control and sensory processing, fundamental to handwriting [63]. The broad gamma activity across these regions suggests a high-level synchronization of cortical activity, facilitating coherent cognitive representations necessary for complex tasks like letter recognition and differentiation [52, 18, 5].
@@
We emphasize that all neurophysiological interpretations are based on correlational analyses of sensor-level EEG signals. Accordingly, observed spectral and spatial differences are interpreted as task-dependent neural associations rather than direct evidence of causal mechanisms or localized neural representations.
@@ VI-B Inter-Trial Coherence in Delta Band During Handwriting Tasks
Studies have indicated the involvement of Delta and Theta bands in motor tasks and cognitive processing [43]. For instance, increased ITC in these bands has been associated with task switching and inhibitory control in BCI paradigms [4]. Similar patterns of coherence in Delta-Theta ranges have been observed during motor performance improvements in stroke rehabilitation, reflecting neural reorganization and recovery [27]. In addition, heightened Theta synchronisation has also been shown to correlate with navigation workload during physical navigation, supporting the role of low-frequency oscillations in coordinating cognitive–motor demands in ecologically valid settings [17, 16]. The periodic auditory stimulation research shows a synchronization peak at 2 Hz in the Delta band, linking sensory-motor time coupling, which may parallel the neural dynamics observed during the rhythmic movements in handwriting [74].
@@ VII Conclusions
This study introduces a CNS that leverages the advanced capabilities of GenAI to enhance spell-based neural language decoding tasks. Our approach is distinct in integrating a CNN with a curriculum-driven LLM, promoting an innovative hybrid method in the domain of BCIs. The framework's effectiveness is demonstrated through its application to EEG-based handwriting of all 26 letters, a novel endeavor in the field. The CNS framework notably achieves exemplary top-k accuracy across all subjects, underscoring the robustness of the EEG encoding model. Furthermore, our curriculum supervised fine-tuning method significantly advances the state of the art by enabling the LLM to effectively learn subject-specific letter transition patterns. This methodological innovation not only enhances sentence synthesis quality but also establishes a system-level strategy for compensating low-SNR neural decoding through language-level error correction. This work addresses a translational feasibility question by demonstrating how full-alphabet neural spelling and curriculum-adapted generative language models can be integrated within non-invasive EEG constraints. By seamlessly merging non-invasive EEG with GenAI, this study advances the methodological foundation of neural spelling systems and sets the stage for future investigations into more constrained paradigms, including online operation and reduced-movement or imagery-based settings.
@@ VIII Limitations
The present study has several limitations. First, all experiments are conducted in an offline decoding setting, and system performance under real-time operation remains to be validated. Second, the current framework primarily relies on within-subject training, which limits cross-subject generalizability and scalability. Third, the dataset size per subject is relatively limited, constraining the training of larger or more expressive neural encoders.
In addition, the proposed neural handwriting paradigm involves overt motor execution, which restricts direct applicability to individuals with severe motor impairments. We therefore position this paradigm as a high-SNR setting for studying full-alphabet neural spelling and language-level correction.
IX Future Work
Future work will extend the proposed framework toward realistic online BCI deployment. Low-latency neural decoding can be achieved using lightweight convolutional models, while language-level correction can operate asynchronously, enabling responsive closed-loop interaction. Future studies will focus on end-to-end online validation, latency analysis, and long-term stability under continuous use, as well as extending the paradigm to reduced-movement or imagery-based settings to improve accessibility. In terms of scalability, future work will prioritize reduced-calibration and few-shot adaptation strategies rather than fully subject-independent decoding, which remains challenging for non-invasive EEG due to substantial inter-subject variability. Future work will also investigate constrained and confidence-aware generative decoding strategies to better balance expressive language reconstruction with reliability and user intent preservation in neural spelling systems. Future work will explicitly address safety and hallucination risks in language-level decoding by incorporating confidence-aware constraints, task-specific vocabularies, and optional human-in-the-loop verification to ensure reliable and intent-preserving communication in clinical settings. Finally, future work will systematically examine behavioral and task-related factors underlying inter-subject variability, including handwriting kinematics (e.g., speed and trajectory complexity) and effective trial counts after artifact rejection, to better contextualize individual differences in decoding performance.